# Kummer-Artin-Schreier-Witt Theory

### v2

@@ Abstract
We prove that the Artin-Schreier-Witt isogeny in characteristic $p>0$ lifts to characteristic 00.
@@
keywords:
Ramification theory, non-archimedean geometry, Kummer-Artin-Schreier-Witt theory.
@@ 1 Introduction
Our motivation is the lifting problem for Galois covers of curves, which asks: given a finite group $G$ and a $G$-cover $\phi$ of a curve in characteristic $p$, does $\phi$ lift to characteristic 00? It is important to note that, in general, the answer is NO. As with many problems in Galois theory, it is natural to begin by investigating the case where $G$ is abelian. There have been significant breakthroughs recently; for instance, a joint effort by Obus, Wewers, and Pop has demonstrated that all cyclic covers of (smooth projective) curves lift [8, 9].
Recall that every abelian cover can be realized as the pullback of a suitable isogeny of algebraic groups [12]. Therefore, to study the lifting problem for abelian covers, one can instead explore the lift of cyclic isogenies from characteristic $p$ to characteristic 00. A classical argument allows us to further restrict to the case where $G$ is cyclic of order $p^{s}$ and the isogeny Weierstrass-$\wp$ of the ring $W_{s}$ of Witt vectors of length $s$, which gives the following exact sequence:
$0\xrightarrow{}\mathbb{Z}/p^{s}\xrightarrow{}W_{s}\xrightarrow{\wp}W_{s}\rightarrow 0,$ (1)
of group schemes over $\overline{k}$, where ${\rm char}(k)=p$. Sekiguchi and Suwa claimed that (1) lifts.
Theorem 1.1 ((c.f. [11, Theorem 8.1])).
For each positive integer $s$, there exists a flat group scheme $\mathcal{W}_{s}$ over $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ that fits into the exact sequence
$0\rightarrow\mathbb{Z}/p^{s}\rightarrow\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}:=\mathcal{W}_{s}/(\mathbb{Z}/p^{s})\rightarrow 0$ (2)
@@
$0\xrightarrow{}\mu_{p^{s}}\xrightarrow{}\mathbb{G}_{m}\xrightarrow{(\cdot)^{p^{s}}}\mathbb{G}_{m}\xrightarrow[]{}0,$ (3)
over $\overline{K}$, where $K=\operatorname{Frac}(R)$. Its special fiber is isomorphic to the Artin-Schreier-Witt sequence (1) over $\overline{\mathbb{F}}_{p}$. Furthermore, $\mathcal{W}_{s}$, as a scheme, is the spectrum of
@@
for explicitly determined polynomials $E_{2},\ldots,E_{s}$ and $\lambda:=\zeta_{p}-1$. Also, $\mathcal{V}_{s}$ is the spectrum of
@@
for explicitly determined polynomials $G_{2},\ldots,G_{s}$.
One immediate implication is the following result.
Theorem 1.2 ((c.f. [11, Theorem 3.8])).
If $B$ is a flat local $R$-algebra, then any unramified $\mathbb{Z}/p^{s}$-cover $\operatorname{Spec}C\rightarrow\operatorname{Spec}B$ is given by a Cartesian diagram of the form
${\operatorname{Spec}C}$${\mathcal{W}_{s}}$${\operatorname{Spec}B}$${\mathcal{V}_{s}}$
However, the paper by Sekiguchi and Suwa remains unpublished. The preprint is 90 pages long and quite challenging for us to read (see [6] for a summary instead). Around the same time, Matsuda developed his own version of Kummer-Artin-Schreier-Witt theory from the perspective of $p$-adic differential equations [4]. More precisely, let $\mathbb{K}$ be a henselian local field of mixed characteristic with imperfect residue field $\kappa$. Matsuda established that the canonical restriction morphism $\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\rightarrow\operatorname{H}^{1}(\mathbb{K},\mathbb{Z}/p^{s})$ is induced by the morphism
$\begin{split}W_{s}(\mathcal{R}^{\operatorname{int}})&\xrightarrow[]{\Psi_{s}}(\mathcal{R}^{\operatorname{int}})^{\times},\\ \underline{X}&\mapsto E_{s,p}(p^{s}\underline{X}),\end{split}$ (6)
where $\mathcal{R}^{\operatorname{int}}$ is a henselian DVR over $\mathbb{K}$, and $E_{s,p}((X_{0},\ldots,X_{s-1}))\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ is a deformation of the Artin-Hasse exponential function.
Specifically, the morphism (6) allows one to study a cyclic extension in mixed characteristic defined by a Kummer class, such as with $E_{s,p}(p^{s}\underline{X})$, by examining the associated Witt vector $\underline{X}$. This approach often simplifies the analysis, as evidenced by many results on Galois extensions in equal-characteristic $p$ obtained through the study of Witt vectors [5, 1].
In this manuscript, we prove the above results using the robust framework developed by Matsuda. More precisely, to recover Theorem 1.1, we apply an algebraization process to (6) to obtain a polynomial $E_{s}(\underline{X})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-1}]$, which is a suitable truncation of $E_{s,p}(p^{s}\underline{X})$. We then show that, for $R=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, the cover
$Z_{s}^{p^{s}}=E_{s}((X_{0},\ldots,X_{s-1}))$
of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ gives (2) after removing the ramification divisor.
@@
Our polynomials $E_{i}$ and $G_{i}$ in (4) and (5) are different from those used by Sekiguchi and Suwa in their work. In fact, these polynomials are not unique, as they are approximations of power series arising from Matsuda theory within a certain degree of accuracy.
Remark 1.4.
We believe that Matsuda’s theory has potential applications far beyond the scope of this paper. For instance, expanding the domain and range of (6) may enable us to address inseparable extensions. This will be explored in our future work on the refined Swan conductor introduced in [2], an important invariant that measure how far a wildly ramified cover is from having separable reduction.
@@ 1.1 Outline
In Section 2, we recall Kummer theory and Artin-Schreier-Witt theory, which classify tamely ramified and wildly ramified cyclic extensions, respectively. In the same section, we restate Matsuda’s theory, which distinguishes our method from that used by Sekiguchi and Suwa. Section 3 defines the cover at the level of field extensions. More precisely, the section constructs the Kummer class of the extension, which corresponds to the $G_{i}$’s in (5), obtained by truncating certain power series arising from Matsuda’s theory. Section 4 derives the $E_{i}$’s in (4) using the previously defined Kummer extension. Section 5 demonstrates that our constructions yield a cyclic étale cover of an affine scheme with good reduction. Section 6 further proves that the cover is an isogeny, establishing Theorem 1.1. Finally, Section 7 provides the proof of Theorem 1.2.
@@ 1.2 Notations
The letter $K$ will always denote a field of mixed characteristic $(0,p)$ with uniformizer $\pi$ and a discrete valuation $\nu:K^{\times}\rightarrow\mathbb{Q}$, normalized so that $\nu(p)=1$. The ring of integers of $K$ will be denoted by $R$, and its residue field $k=R/(\pi)$ is perfect. An example to keep in mind is $K=\mathbb{Q}[\zeta_{p^{s}}]$, with $\pi=\zeta_{p^{s}}-1$, which has valuation $\nu(\pi)=\frac{1}{p^{s-1}(p-1)}$. Set $\lambda=\zeta_{p}-1$, a constant that will appear frequently.
We fix an algebraic closure $\overline{K}$ of $K$. The symbol $\mathbb{K}$ usually denotes a Laurent series over $K$. The valuation $\nu$ extends to the Gauss valuation on $\mathbb{K}$. For instance, if $\mathbb{K}=K\mathopen{(\!(}X\mathclose{)\!)}$ and $f(X)=\sum_{i\geq n}a_{i}X^{i}\in\mathbb{K}$, then $\nu(f)=\inf_{i\geq n}\{\nu(a_{i})\}$.
1.3 Acknowledgements
We thank Kiran Kedlaya, Ariane Mézard, Andrew Obus, Frans Oort, Matthieu Romagny, and Dajano Tossici for their valuable discussions. The first author expresses gratitude to the National Center for Theoretical Sciences at National Taiwan University for providing excellent working conditions.
@@ 2 Background
2.1 Artin-Schreier-Witt theory
Suppose $k$ is a field of characteristic $p$. Then we have the short exact sequence:
$\begin{split}0\xrightarrow{}\mathbb{Z}/p^{s}\cong W_{s}(\mathbb{F}_{p})\xrightarrow{}W_{s}&\xrightarrow{\wp}W_{s}\rightarrow 0,\\ \underline{y}=(y_{0},\ldots,y_{s-1})&\mapsto\underline{y}-\sigma(\underline{y}),\end{split}$ (7)
of group schemes over $\overline{k}$, where $\sigma$ is the Frobenius operator on Witt vectors. This induces the identification
@@ 2.1 Artin-Schreier-Witt theory
One may hence view $\wp:W_{s}\rightarrow W_{s}$ in (7) as a $\mathbb{Z}/p^{s}$-extension of $\mathcal{K}:=\overline{k}(x_{0},\ldots,x_{s-1})$ given by the Witt vector $(x_{0},\ldots,x_{s-1})$. Furthermore, if Weierstrass-$\wp$ is associated with a character $\overline{\chi}\in\operatorname{H}^{1}(\mathcal{K},\mathbb{Z}/p^{s})$, then $\overline{\chi}^{\otimes p^{s-j}}$ is the $\mathbb{Z}/p^{j}$-subextension corresponding to the Witt vector $(x_{0},\ldots,x_{j-1})$.
Additionally, if we write
@@
where $f_{i}(y_{0},\ldots,y_{i-2})\in\mathbb{F}_{p}(y_{0},\ldots,y_{i-2})$ for $i=2,\ldots,s$ and $f_{1}=0$, then $\overline{\chi}$ is defined by a system of Artin-Schreier extensions
@@ 2.2 Kummer Theory
Fix $m\in\mathbb{Z}_{\geq 0}$. Suppose $K$ contains a primitive $m$-th root of unity and has characteristic prime to $m$. By Kummer theory, we have the following isomorphism:
@@
The isomorphism follows from the exact sequence over $\overline{K}$:
$0\rightarrow\mu_{m}\rightarrow\overline{K}^{\times}\xrightarrow{(\cdot)^{m}}\overline{K}^{\times}\rightarrow 0$
Thus, each $\chi\in\operatorname{H}^{1}(K,\mathbb{Z}/p^{s}\mathbb{Z})$ can be associated with a rational function $G$ in $K^{\times}/(K^{\times})^{p^{s}}$. Additionally, we may assume that the function $G$ is of the following form:
$G:=\prod_{i=0}^{s-1}G_{i}^{p^{i}},$
where each $G_{i}\in K\setminus K^{p}$ or $G_{i}=1$. As in Section 2.1, $\chi^{\otimes p^{s-j}}$ corresponds to a character in $\operatorname{H}^{1}(K,\mathbb{Z}/p^{j}\mathbb{Z})$ defined by the Kummer extension:
@@
In fact, $\chi$ can be defined by a system of Kummer extensions of degree $p$:
@@
where $Z_{-1}=1$.
We thus establish the following isomorphism:
$(K^{\times})^{\oplus s}/\phi_{s}\big{(}(K^{\times})^{\oplus s}\big{)}$ $\xrightarrow{\sim}$ $\;K^{\times}/\big{(}K^{\times}\big{)}^{p^{s}}$ $\xrightarrow{\sim}\operatorname{H}^{1}\big{(}K,\mathbb{Z}/p^{s}\mathbb{Z}\big{)}$ (9)
$(a_{1},\ldots,a_{s})$ $\mapsto$ $\;[a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}]$ $\mapsto\mathfrak{K}_{s}\big{(}a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}\big{)},$
@@
$(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ is a group under coordinate-wise multiplication. We say $(a_{1},\ldots\allowbreak\ldots,a_{s})\sim(b_{1},\ldots,b_{s})$ if there exists $(c_{1},\ldots,c_{s})$ such that
@@
This equivalence relation is induced from $K^{\times}/(K^{\times})^{p^{s}}$.
This leads to a Kummer-like sequence:
$0\xrightarrow{}\mu_{p^{s}}\xrightarrow{}(\mathbb{G}_{m})^{s}\xrightarrow{\phi_{s}}(\mathbb{G}_{m})^{s}\xrightarrow[]{}0,$ (10)
where the injection sends $1$ to $(\zeta_{p},\zeta_{p^{2}},\ldots,\zeta_{p^{s}})$.
2.3 Kummer-Artin-Schreier-Witt theory
In this section, we introduce Matsuda’s theory as presented in [3].
2.3.1 Deformed Artin-Hasse exponential function
We first introduce a deformation of the Artin-Hasse exponential function and some of its properties.

### v3

# Kummer–Artin–Schreier–Witt Theory

@@ Abstract
We study the problem of lifting the Artin–Schreier–Witt isogeny from characteristic $p>0$ to characteristic 0, which is central to the lifting problem for Galois covers of algebraic schemes in positive characteristic. We introduce a new technique that associates a Kummer class, representing a tamely ramified cyclic extension, to a Witt vector via Matsuda’s Kummer–Artin–Schreier–Witt theory. This viewpoint leads to an explicit construction of a lift of the isogeny over a concrete base ring. Our results lay the groundwork for further applications, including the study of inseparable extensions and Kato’s refined Swan conductor.
keywords:
Ramification theory, non-archimedean geometry, Kummer–Artin–Schreier–Witt theory.
@@ 1 Introduction
Our motivation is the lifting problem for Galois covers of curves, which asks whether the following holds: given a finite group $G$ and a $G$-cover $\phi$ of a curve in characteristic $p$, does $\phi$ lift to characteristic 0? In general, the answer is negative [17, §1.1]. As with many problems in Galois theory, it is natural to begin with the case where $G$ is abelian. There have been significant recent breakthroughs; for instance, work of Obus–Wewers [18] and Pop [19] shows that all cyclic covers of smooth projective curves lift [19, Theorem 1.1].
Recall that every abelian cover can be realized as the pullback of a suitable isogeny of algebraic groups [23, I, Theorem 4]. Therefore, to study the lifting problem for abelian covers, one may instead consider lifting cyclic isogenies from characteristic $p$ to characteristic 0. A classical argument further reduces to the case where $G$ is cyclic of order $p^{s}$, with isogeny $\wp$ on the ring $W_{s}$ of Witt vectors of length $s$, yielding the following exact sequence:
[…]
of group schemes over an algebraically closed field $\overline{k}$ of characteristic $p$. Sekiguchi and Suwa showed that (1) lifts.
Theorem 1.1 (cf. [21, Theorem 8.1]).
For each positive integer $s$, there exists a flat group scheme $\mathcal{W}_{s}$ over $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ that fits into an exact sequence
$0\longrightarrow\mathbb{Z}/p^{s}\longrightarrow\mathcal{W}_{s}\longrightarrow\mathcal{V}_{s}:=\mathcal{W}_{s}/(\mathbb{Z}/p^{s})\longrightarrow 0$ (2)
@@
[…]
over $\overline{K}$, where $K=\operatorname{Frac}(R)$. Its special fiber is isomorphic to the Artin–Schreier–Witt sequence (1) over $\overline{\mathbb{F}}_{p}$. Furthermore, $\mathcal{W}_{s}$, as a scheme, is the spectrum of
@@
for explicitly determined polynomials $E_{2},\ldots,E_{s}$ with coefficients in $R$, and $\lambda:=\zeta_{p}-1$. Likewise, $\mathcal{V}_{s}$ is the spectrum of
@@
for explicitly determined polynomials $G_{2},\ldots,G_{s}$ with coefficients in $R$.
One immediate implication of Theorem 1.1 is the following result.
Theorem 1.2 (cf. [21, Theorem 3.8]).
If $B$ is a flat local $R$-algebra, then any unramified $\mathbb{Z}/p^{s}$-cover $\operatorname{Spec}C\to\operatorname{Spec}B$ is given by a Cartesian diagram of the form
$\operatorname{Spec}C$$\mathcal{W}_{s}$$\operatorname{Spec}B$$\mathcal{V}_{s}.$}
However, the paper by Sekiguchi and Suwa remains unpublished; see [16] for a summary. Around the same time, Matsuda developed his own version of Kummer–Artin–Schreier–Witt theory from the perspective of $p$-adic differential equations [14]. More precisely, let $\mathbb{K}$ be a Henselian local field of mixed characteristic $(0,p)$ with imperfect residue field $\kappa$. Matsuda established that the canonical restriction morphism $\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\rightarrow\operatorname{H}^{1}(\mathbb{K},\mathbb{Z}/p^{s})$ is induced by the morphism
[…]
where $\mathcal{R}^{\operatorname{int}}$ is a Henselian discrete valuation ring over $\mathbb{K}$ (for more details, see § 2.3.2), and $E_{s,p}((X_{0},\ldots,X_{s-1}))\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ is a deformation of the Artin–Hasse exponential function.
Specifically, the morphism (6) allows one to study a cyclic extension in mixed characteristic defined by a Kummer class (for example, one represented by $E_{s,p}(p^{s}\underline{X})$) by examining the associated Witt vector $\underline{X}$. This approach often simplifies the analysis, as illustrated by many results on Galois extensions in equal characteristic $p$ obtained through the study of Witt vectors [15, 6].
In this paper, we reprove the above results using Matsuda’s framework. More precisely, to recover Theorem 1.1, we apply an algebraization process to (6) to obtain a polynomial $E_{s}(\underline{X})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-1}]$, which is a suitable truncation of $E_{s,p}(p^{s}\underline{X})$. We then show that, for $R=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, the cover
$Z_{s}^{p^{s}}=E_{s}(X_{0},\ldots,X_{s-1})$ (7)
of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ induces (2) after removing the ramification divisor.
@@
Our polynomials $E_{1},\ldots,E_{s}$ and $G_{1},\ldots,G_{s}$ in (4) and (5), defined in Definition 4.9 and Proposition 3.8, differ from those used by Sekiguchi and Suwa. In fact, these polynomials are not unique, as they approximate power series arising from Matsuda’s theory within a certain degree of accuracy.
The key strategy in the proof is to associate a Kummer class, which represents a cyclic extension of degree $p^{s}$ of a valued field $\mathbb{K}$ of mixed characteristic $(0,p)$, with a Witt vector $\underline{X}=(X_{0},\ldots,X_{s-1})\in W_{s}(\mathbb{K})$ using Matsuda’s theory (see Proposition 3.4). We believe that this technique has potential applications beyond the scope of this paper. For instance, enlarging the domain and codomain of (6) may allow us to address inseparable extensions and to study the refined Swan conductor [8], an important invariant measuring how far a wildly ramified extension of a field with imperfect residue field is from having separable reduction.
When the field $\mathbb{K}$ is of equal characteristic $p$ with imperfect residue field, any $\mathbb{Z}/p^{s}$-extension of $\mathbb{K}$ can be represented by a Witt vector $\underline{x}=(x_{0},\ldots,x_{s-1})\in W_{s}(\mathbb{K})$ (see §2.1). Moreover, when $\underline{x}$ satisfies certain conditions, the refined Swan conductor is completely determined by $\underline{x}$ (see [8, Lemma 3.7], [12, Proposition 2.8]). We conjecture that an analogue of this result holds in the mixed characteristic setting, where $\underline{X}$ plays the role of $\underline{x}$. In this paper, we prove the conjecture in the case $s=1$.
Proposition 1.4.
Let $\mathbb{K}$ be a field of mixed characteristic $(0,p)$ with uniformizer $\pi$, residue field $\kappa$, and a valuation $\nu$ normalized by $\nu(p)=1$. Let $e$ be the absolute ramification index. Let $(X_{0})\in W_{1}(\mathbb{K})$ denote a Witt vector associated with a $\mathbb{Z}/p$-extension of $\mathbb{K}$. For $t\in\mathbb{R}$, set
$\mathfrak{p}_{\mathbb{K}}^{\,t}:=\{\,x\in\mathbb{K}\mid\nu(x)\geq t\,\}.$
Assume that
$\frac{p}{p-1}<\nu(X_{0})<0,$
and define
$x_{0}:=X_{0}\,\pi^{-e\nu(X_{0})}.$
Suppose that the reduction $\overline{x}_{0}$ is not contained in $\kappa^{p}$. Then the refined Swan conductor of the associated character $\mathscr{F}$ is
$\mathrm{rsw}(\mathscr{F})=-dX_{0}\otimes 1=-\pi^{e\nu(X_{0})}\otimes d\overline{x}_{0}\in\mathfrak{p}_{\mathbb{K}}^{-{\mathrm{sw}}(\mathscr{F})}\otimes_{\mathcal{O}_{\mathbb{K}}}\Omega^{1}_{\kappa},$
where $\mathrm{sw}(\mathscr{F})=-\nu(X_{0})$ is the Swan conductor of $\mathscr{F}$.
@@ 1.1 Outline
In Section 2, we review Kummer theory and Artin–Schreier–Witt theory, which classify tamely and wildly ramified cyclic extensions, respectively. We also recall Matsuda’s Kummer–Artin–Schreier–Witt theory and explain how it leads to our explicit construction.
Section 3 constructs, at the level of field extensions, the cover of $\mathbb{P}^{s}_{R}$ defined by (7). Concretely, we build the relevant Kummer class using the elements $G_{1},\ldots,G_{s}$ from (5), obtained by truncating the power series arising from Matsuda’s theory. From this Kummer extension, Section 4 derives the polynomials $E_{1},\ldots,E_{s}$ in (4).
In Section 5, we show that these constructions yield a cyclic étale cover of an affine scheme with good reduction. Section 6 studies the underlying algebraic group structures via Witt vector operations; using Hopf algebras and descent, we prove that the cover is an isogeny, establishing Theorem 1.1. Section 7 proves Theorem 1.2. Finally, Section 8 discusses the refined Swan conductor for cyclic characters and proves Proposition 1.4.
@@ 1.2 Notations
The letter $K$ will always denote a field of mixed characteristic $(0,p)$ with uniformizer $\pi$ and a discrete valuation $\nu:K^{\times}\to\mathbb{Q}$, normalized so that $\nu(p)=1$. We set $e:=\nu(\pi)$ for the absolute ramification index. For any $u\in\mathcal{O}_{K}$, we denote by $\overline{u}$ its image in the residue field. The ring of integers of $K$ will be written $R$, and its residue field $k=R/(\pi)$ is assumed to be perfect. An example to keep in mind is $K=\mathbb{Q}(\zeta_{p^{s}})$ with $\pi=\zeta_{p^{s}}-1$, for which $\nu(\pi)=\tfrac{1}{p^{s-1}(p-1)}$. We also set $\lambda:=\zeta_{p}-1$, a constant that will appear frequently.
We fix an algebraic closure $\overline{K}$ of $K$. We write $\mathbb{K}$ for a Henselian local subfield of the Laurent series field $K\!\mathopen{((}X\mathclose{))}$; for instance, one may take $K\!\mathopen{((}X\mathclose{))}^{\mathrm{bd}}$, the subfield consisting of Laurent series with bounded coefficients. The valuation $\nu$ extends naturally to the Gauss valuation on $\mathbb{K}$. For example, if $f(X)=\sum_{i\geq n}a_{i}X^{i}\in\mathbb{K}$, then
$\nu(f)=\inf_{i\geq n}\{\nu(a_{i})\}.$
Acknowledgements
We thank Kiran Kedlaya, Ariane Mézard, Andrew Obus, Frans Oort, Matthieu Romagny, and Dajano Tossici for valuable discussions. We are especially grateful to Adrian Vasiu for carefully reading an earlier version of this manuscript.
@@ 2 Background
In this section, we briefly review the classical theories that form the foundation of our work: Artin–Schreier–Witt theory, which governs wildly ramified cyclic extensions in characteristic $p$; Kummer theory, which describes tamely ramified cyclic extensions; and Matsuda’s Kummer–Artin–Schreier–Witt theory, which provides a unifying framework in mixed characteristic $(0,p)$.
2.1 Artin–Schreier–Witt theory
Suppose $k$ is a field of characteristic $p$. Then we have the short exact sequence
[…]
of group schemes over $\overline{k}$, where $\sigma:(y_{0},\ldots,y_{s-1})\mapsto(y_{0}^{p},\ldots,y_{s-1}^{p})$ is the Frobenius operator on Witt vectors. By Hilbert’s Theorem 90, this induces the identification
@@ 2.1 Artin–Schreier–Witt theory
We denote by $\mathfrak{K}_{s}(\underline{x})$ the character in $\operatorname{H}^{1}(k,\mathbb{Z}/p^{s})$ associated with the Witt vector $\underline{x}\in W_{s}(k)$.
One may view $\wp:W_{s}\to W_{s}$ in (8) as defining a $\mathbb{Z}/p^{s}$-extension of the function field $\mathcal{K}:=\overline{k}(x_{0},\ldots,x_{s-1})$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$. Furthermore, if $\wp$ corresponds to a character $\overline{\chi}\in\operatorname{H}^{1}(\mathcal{K},\mathbb{Z}/p^{s})$, then $\overline{\chi}^{\otimes p^{s-j}}$ is the $\mathbb{Z}/p^{j}$-subextension corresponding to the Witt vector $(x_{0},\ldots,x_{j-1})$.
Moreover, if we write
@@
where $f_{i}(y_{0},\ldots,y_{i-2})\in\mathbb{F}_{p}(y_{0},\ldots,y_{i-2})$ for $i=2,\ldots,s$ and $f_{1}=0$, then $\overline{\chi}$ is defined by the system of Artin–Schreier equations
@@ 2.2 Kummer Theory
Fix $m\in\mathbb{Z}_{>0}$. Suppose $K$ contains a primitive $m$-th root of unity and has characteristic coprime to $m$. By Kummer theory, we have the isomorphism
@@
This follows from the Kummer exact sequence over $\overline{K}$:
[…]
We now specialize to the case $m=p^{s}$. In particular, each $\chi\in\operatorname{H}^{1}(K,\mathbb{Z}/p^{s})$ corresponds to an element $G\in K^{\times}/(K^{\times})^{p^{s}}$.
Proposition 2.1.
Every class $G]\in K^{\times}/(K^{\times})^{p^{s}}$ has a representative of the form
$G\equiv\prod_{i=0}^{s-1}G_{i}^{p^{i}}\quad\big(\mathrm{mod}\ (K^{\times})^{p^{s}}\big),$
with each $G_{i}\in K^{\times}\setminus(K^{\times})^{p}$ or $G_{i}=1$.
Proof.
Choose a set $S\subset K^{\times}$ of representatives for $K^{\times}/(K^{\times})^{p}$ such that $1\in S$ and, for every $x\in S\setminus\{1\}$, we have $x\notin(K^{\times})^{p}$.
If $s=1$, then any $G$ is congruent modulo $(K^{\times})^{p}$ to some $G_{0}\in S$, so $G\equiv G_{0}^{p^{0}}$.
Now suppose the statement holds for $s-1\geq 1$, i.e., $s\geq 2$. Take $g\in K^{\times}$ representing $G]\in K^{\times}/(K^{\times})^{p^{s}}$. Choose $G_{0}\in S$ such that $g\in G_{0}(K^{\times})^{p}$, i.e.
$g=G_{0}a^{p},\qquad a\in K^{\times}.$
By the induction hypothesis, there exist $H_{0},\dots,H_{s-2}\in S$ with
$a\equiv\prod_{i=0}^{s-2}H_{i}^{p^{i}}\quad\big(\mathrm{mod}\ (K^{\times})^{p^{s-1}}\big).$
Raising this congruence to the $p$-th power and multiplying by $G_{0}$, we obtain
$g\equiv G_{0}\left(\prod_{i=0}^{s-2}H_{i}^{p^{i}}\right)^{p}=G_{0}\prod_{i=0}^{s-2}H_{i}^{p^{i+1}}=\prod_{i=0}^{s-1}G_{i}^{p^{i}}\quad\big(\mathrm{mod}\ (K^{\times})^{p^{s}}\big),$
where $G_{0}$ is as chosen and $G_{i}:=H_{i-1}$ for $i\geq 1$. Each $G_{i}\in S$, hence either $1$ or not a $p$-th power. This completes the induction. ∎
As before, we write $\chi=\mathfrak{K}_{s}(G)$. As in Section 2.1, $\chi^{\otimes p^{s-j}}$ corresponds to a character in $\operatorname{H}^{1}(K,\mathbb{Z}/p^{j})$ defined by the Kummer extension
@@
In fact, $\chi$ can be described by a system of successive Kummer extensions of degree $p$:
@@
with $Z_{-1}=1$.
We thus obtain canonical isomorphisms:
[…]
$(a_{1},\ldots,a_{s})$ $\longmapsto$ $\;[a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}]$ $\longmapsto\mathfrak{K}_{s}\big(a_{1}a_{2}^{p}\cdots a_{s}^{p^{s-1}}\big),$
@@
The quotient $(K^{\times})^{\oplus s}/\phi_{s}((K^{\times})^{\oplus s})$ is a group under coordinatewise multiplication. We declare
$(a_{1},\ldots,a_{s})\sim(b_{1},\ldots,b_{s})$
if there exists $(c_{1},\ldots,c_{s})$ such that
@@
This is the equivalence relation induced by $K^{\times}/(K^{\times})^{p^{s}}$.
This yields a Kummer-type exact sequence of group schemes:
[…]
where $\iota$ is determined by
$\iota(\zeta_{p^{s}})=(\zeta_{p},\zeta_{p^{2}},\ldots,\zeta_{p^{s}}).$
2.3 Kummer–Artin–Schreier–Witt theory
We review Matsuda’s theory [14] and follow the notation of [10, §9].
2.3.1 Deformed Artin–Hasse exponential function
We introduce a deformation of the Artin–Hasse exponential function and record basic properties.